OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = x + 8 x= 4over3 y = x + 3 x=4over3 y = −x + 8 -x= -3over4 y = −x + 3 -x= -3over4

jimthompson5910 (jim_thompson5910):

4x − 3y = −12 -3y = -4x - 12 y = 4/3x + 4 So the slope is 4/3

jimthompson5910 (jim_thompson5910):

Any parallel line will also have a slope of 4/3

jimthompson5910 (jim_thompson5910):

y = mx+b y = (4/3)x + b 4 = (4/3)(-3) + b 4 = -12/3 + b 4 = -4 + b 4+4 = b 8 = b b = 8 So the equation of the parallel line is \(\Large y = \frac{4}{3} x + 8\)